OpenStudy (laronjaxon):

explain how you would use the quadratic formula to solve x2 + 7x = –3.

OpenStudy (anonymous):

x²+7x = -3 x²+7x+3 = 0 a = 1, b=7, c=3 x = [-b ± sqrt(b²-4ac)] / 2a = [-7 ± sqrt( (-7)² - 4(1)(3) )] / 2(1) = (-7 ± sqrt( 37 )) / 2

OpenStudy (anonymous):

im not sure if thats right.. but thats my answer..

OpenStudy (laronjaxon):

lol thank you

OpenStudy (lalaly):

correction ,, its sqrt(7^2-4(1)(3)